# ACT-R: Adaptive Camera Trajectories for 3D Reconstruction from Single Image

### v1

@@ 4.2 Quantitative Comparisons
Since the reconstructed 3D meshes from different methods exhibit different canonical poses, we first normalize the output mesh and provide an initial transform to align it with the GT mesh, then optionally used Iterative Closest Point (ICP) to further tune the mesh poses, whichever leads to better quantitative measures for the results.
We report quantitative comparison results for all 1,030 objects from GSO in Table 1. Our method demonstrates clear advantages over the other methods across all metrics, especially in F1, a critical and meaningful metric for 3D shape generation/reconstruction [57]. The improvements on 2D metrics indicate that our method can generate objects that are aligned with the GT in pixel space. Overall, our method exhibits robustness in accurately modeling object geometry while preserving visual realism in image-space projections. Its superior performance on GSO further underscores its strong generalizability across diverse object categories, showcasing its capability in handling a wide range of shapes, sizes, and appearances effectively.
@@
$\text{Ours}_{\text{lrm}}$ 4.51 3.41 16.2 17.4 8.49 14.0 1.05
@@ 4.3 Visualizations of Trajectories
Fig. 6 shows our predicted slice images and trajectories. An naive view planning would be to negate on the initial elevation angle for an even span, i.e., raising the camera for negative initial elevation angles and lowering it for positive ones. In contrast, we based our trajectory prediction on a deeper understanding of object structures and occlusions.
For the bundt pan, even with a positive elevation for the input view, we still raise the camera to reveal the inner tube. Although slice3D [64] failed to predict the inner tube in the sliced image, it still suggests that inner regions require additional attention. For the tape example, we can see the camera highest elevation is lower that the pan example, since it has already fully observed the hole in the middle. This proves that our method is adaptive to each single object. For the sneaker example, our trajectory can better observe the pair of wings so that they will not occlude with each other in most of the frames. For the sofa, we lower the camera to fully capture its bottom. For the Jenga blocks, we planned a trajectory that can better observe the concave areas.
@@ 4.4 Ablation Studies
Robustness to camera pose estimation. Our method relies on estimating camera parameters and the bounding box. Since our trajectory planning works on a coarse 3D representation (blocks), it shows robustness against estimation errors. More details on how camera estimation affects our view planning can be found in the Supplementary.
@@ 5 Conclusion, Limitation, and Future Work
We propose adaptive view planning for synthesizing novel views from a single image. Our key insight is that slice images from Slice3D [64] can effectively reveal occluded structures from the input view and guide the camera’s movement. By leveraging the capabilities of video generation models, our planned trajectory generates novel views maximises the visibility of occluded structures while maintaining multi-view consistency and the overall frame budget. Interestingly, this reveals how multi-slice and multi-view can work together to complement each other, where the former provides geometric insights on where the self-occlusion may be present while the later offers signals on the visible geometries. Combining the two, potential invisible regions are better exposed for improved 3D reconstruction.
Most limitations in our current implementation are inherited from SV3D [60], e.g., limited camera DoFs as only azimuth and elevation are changeable, and quality of the generated videos. Nonetheless, our adaptive view planning is applicable to other video generation models such as ViewCrafter [73], VD3D [1], and CameraCtrl [14] with even higher camera DoFs. ACT-R’s performance is also limited by the accuracy of the estimated semantic differences between slice and input images. It is certainly possible to explore alternative occlusion-aware reasoning for view planning, or enlarge the scale of 3D data exposed to Slice3D — the version we employed for trajectory planning was trained on only 5% of the 880K Objaverse 3D dataset.
At last, an intriguing discussion rests on a conundrum over adaptive vs. fixed view syntheses. While we firmly believe in the merits of adaptive view trajectories for occlusion revelation, view synthesis from such trajectories (e.g., as supported by SV3D [60]) is more difficult compared to one that only works with fixed cameras, e.g., Zero123++ [50]. Even when trained on the same 3D dataset, SV3D is clearly outdone by Zero123++ in terms of the quality of the synthesized multi-view images. As a result, our method, which feeds adaptive trajectories to SV3D, cannot quantitatively beat InstantMesh [26], which employs Zero123++, across most metrics. Qualitatively however, we consistently observe that our method can outperform InstantMesh when the 3D objects have significant concavities or occlusions; see the two bowls in the fourth blue column in Fig. 4. We are motivated to resolve the above discrepancy by boosting the performance of camera-adaptive view synthesis.
@@
6 Estimation of Camera Poses and Object Bounding Boxes
@@
For the bounding boxes pediction, we directly use MSE:
@@
7 Robustness Against Camera Pose Errors
@@
Figure 9: Robustness to camera pose prediction errors. The left and right show our predicted camera trajectories using predicted and GT camera poses, respectively. The numbers in brackets show the elevation changes in each segment of the orbit. ”Vis. Score” denotes the visibility score defined in Equation 2.
8 Visibility Set Calculation
@@
To calculate the visibility set $\psi(\pi(t))$ in Equation (2), we perform two key checks for each difference block: (1) a field of view check and (2) an occlusion check. The field of view check ensures blocks lie within the camera’s viewing angle (33.6°), computed as the angle between the camera direction and the vector from camera to block center. The occlusion check determines if any other block intersects the line segment from camera to the target block’s center using a ray-box intersection algorithm. A block is considered visible only if it passes both checks. This visibility function allows us to evaluate each candidate trajectory by summing the weights of all visible blocks across all camera positions, leading to the selection of the optimal trajectory $\pi^{*}$ that maximizes the visibility of important difference regions.
9 Calculation of Random Trajectory
@@
Following the same set up as our adaptive trajectories, we apply uniform azimuth angle change of $18^{\circ}$ and divide the camera trajectory orbit into four segments. Instead of setting a constant elevation increment stepsize within each segment, we randomly sample an incrementation step with $[-5^{\circ},5^{\circ}]$ for each time of elevation change. To ensure the final trajectory forms a close loop, we apply the same mirrong and negating mechanism between segments.
10 More Visual Results
Using object slices as priors, our adaptive trajectory method deliberately captures regions occluded from the input view. As shown in Fig. 10, our model demonstrates consistent performance across various challenging cases: hallucinating the right Jenga’s crease (row 1), reconstructing solid bowl bottoms with the right cavity (rows 2 and 3), revealing hidden nose and tails (row 4,5,7), and correctly modeling the parallel wings (row 6). These results demonstrate that our minimally-trained adaptive trajectory approach yields substantial improvements in reconstruction quality.
Figure 10: Additional visual comparison between single-view 3D reconstruction methods on GSO dataset. Different colors are used to distinguish meshes generated by each method. Our slice-guided approach demonstrates superior reconstruction of occluded regions, including bowl bottoms, hidden wings and extended features like raccoon tails and elephant nose. Please zoom in for detailed comparison of the reconstruction quality.

### v2

@@ 4.2 Quantitative Comparisons
Since the reconstructed 3D meshes from different methods exhibit different poses, we first normalize the output mesh and provide an initial transform to align it with the GT mesh, then optionally used Iterative Closest Point (ICP) to further tune the mesh poses, whichever leads to better quantitative measures for the results.
We report quantitative comparison results for all 1,030 objects from the GSO dataset in Table 1. Our method demonstrates clear advantages over the other methods across all metrics, especially in F1 [75]. The improvements on 2D metrics indicate that our method can generate objects that are aligned with the GT in pixel space. Overall, our method exhibits robustness in accurately modeling object geometry while preserving visual realism in image-space projections. Its superior performance on GSO further underscores its strong generalizability across diverse object categories, showcasing its capability in handling a wide range of shapes, sizes, and appearances effectively.
@@
$\text{Ours}_{\text{IM}}$ 4.51 3.41 16.2 17.4 8.49 14.0 1.05
@@ 4.3 Visualizations of Trajectories
Fig. 5 shows our predicted slice images and trajectories. An naive view planning would negate the initial elevation angle for an even span, raising the camera for negative angles and lowering it for positive ones. In contrast, we base our trajectory prediction on a deeper understanding of object structures and occlusions.
For the Bundt cake pan, even with a positive elevation for the input view, we still raise the camera to reveal the inner tube. Although slice3D [3] failed to predict the inner tube in the sliced image, it still suggests that inner regions require additional attention. For the tape example, we can see the camera highest elevation is lower that the pan example, since it has already fully observed the hole in the middle. This proves that our method is adaptive to each single object. For the sneaker example, our trajectory can better observe the pair of wings so that they will not occlude with each other in most of the frames. For the sofa, we lower the camera to fully capture its bottom. For the Jenga blocks, we planned a trajectory that can better observe the concave areas.
@@ 4.4 Ablation Studies
Robustness to camera pose estimation. Since our trajectory planning operates on a coarse 3D representation, i.e., the blocks, it is robust against errors from estimating the camera parameters and the bounding boxes. More details on how camera estimation affects our view planning can be found in the Supplementary.
@@ 5 Conclusion, Limitation, and Future Work
We propose adaptive view planning for synthesizing novel views from a single image. Our key insight is that slice images from Slice3D [3] can effectively reveal occluded structures from the input view and guide the camera’s movement. By leveraging the capabilities of modern video generation models, the generated novel views along our planned trajectory tend to improve the visibility of occluded structures while maintaining multi-view consistency and the overall frame budget. Interestingly, this reveals how multi-slice and multi-view can work together to complement each other, where the former provides geometric insights on where the self-occlusion may be present while the later offers signals on the visible geometries. Combining the two, potential invisible regions are better exposed for improved single-view 3D reconstruction. In the end, our method ACT-R has been shown to outperform state-of-the-art reconstruction alternatives which take the single-to-multi view route without resorting to direct 3D supervision.
Limitations. Most limitations in our current implementation are inherited from SV3D [1], e.g., limited camera DoFs as only azimuth and elevation are changeable, and quality of the generated videos. Nonetheless, our view planning is applicable to other video generation models such as ViewCrafter [49], VD3D [78], and CameraCtrl [47] with even higher camera DoFs. ACT-R’s performance is also limited by the accuracy of the estimated semantic differences between slice and input images. It is certainly possible to explore alternative occlusion-aware reasoning for view planning, or enlarge the scale of 3D data exposed to Slice3D — the version we employed for trajectory planning was trained on only 5% of the 880K Objaverse 3D dataset.
Adaptive vs. fixed camera trajectories. While we firmly believe in the merits of adaptive view trajectories for occlusion revelation, view synthesis from such trajectories (e.g., via SV3D [1]) is more difficult compared to one that only works with fixed cameras, e.g., Zero123++ [18]. Even when trained on the same 3D dataset, SV3D is clearly outdone by Zero123++ in terms of the quality of the synthesized multi-view images. As a result, our method, which feeds adaptive trajectories to SV3D, cannot quantitatively beat InstantMesh [21], which employs Zero123++, across most metrics. Qualitatively however, we consistently observe that our method can outperform InstantMesh when the 3D objects have significant concavities or occlusions; see the two bowls in the fourth blue column in Fig. 9. We are motivated to resolve the above discrepancy by improving camera-adaptive view synthesis.
@@
LRMs with direct 3D supervision. The most successful single-view reconstruction models of late have predominantly been LRMs [6, 10, 9] with direct image-to-3D and 3D supervision using large-scale 3D datasets. The strong 3D priors learned by these models appears to be diminishing the importance of multi-view synthesis. As we show in the supplementary material, multi-view Trellis does not outperform its single-view counterpart. This is also due in part to the blurriness of the synthesized multi-view images, regardless of the camera trajectories. That being said, hallucinations still abound either due to severe occlusions in the input view (see Fig. 3) or erroneous priors (e.g., symmetries by CLAY [6]), while potential overfitting to 3D training data remains a valid concern.
In future work, a more in-depth investigation into how to best integrate adaptive view planning into direct image-to-3D LRMs is worth conducting. We would also like to explore other applications of ACT, e.g., for robotics and autoscanning [79].
Figure 10: Additional qualitative visual comparisons between single-view 3D reconstruction methods on the GSO dataset. Please zoom in for a closer inspection. Meshes reconstructed with NeUS and LRM are in yellow and blue blocks, respectively. Our slice-guided approach demonstrates superior reconstruction of occluded regions, including bowl bottoms, hidden wings and extended features like raccoon tails and elephant trunk. Please zoom in for detailed comparison of the reconstruction quality.
@@
Appendix A Estimation of Camera Poses and Object Bounding Boxes
@@
For the bounding boxes prediction, we directly use MSE:
@@
Appendix B Robustness Against Camera Pose Errors
@@
Figure 11: Robustness to camera pose prediction errors. The left and right show our predicted camera trajectories using predicted and GT camera poses, respectively. The numbers in brackets show the elevation changes in each segment of the orbit. ”Vis. Score” denotes the visibility score defined in Equation (1) from the main paper.
Appendix C Visibility Set Calculation
@@
To calculate the visibility set $\psi(\pi(t))$ in Equation (1) from the main paper, we perform two key checks for each difference block: (1) a field of view check and (2) an occlusion check. The field of view check ensures blocks lie within the camera’s viewing angle (33.6°), computed as the angle between the camera direction and the vector from camera to block center. The occlusion check determines if any other block intersects the line segment from camera to the target block’s center using a ray-box intersection algorithm. A block is considered visible only if it passes both checks. This visibility function allows us to evaluate each candidate trajectory by summing the weights of all visible blocks across all camera positions, leading to the selection of the optimal trajectory $\pi^{*}$ that maximizes the visibility of important difference regions.
Appendix D Calculation of Random Trajectory
@@
Following the same set up as our adaptive trajectories, we apply uniform azimuth angle change of $18^{\circ}$ and divide the camera trajectory orbit into four segments. Instead of setting a constant elevation increment stepsize within each segment, we randomly sample an incrementation step with $[-5^{\circ},5^{\circ}]$ for each time of elevation change. To ensure the final trajectory forms a close loop, we apply the same mirroring and negating mechanism between segments.
Appendix E Alternative Feature Encoders
In our primary experiments, we utilized VGG16 [4] as the feature extractor to quantify semantic differences between slice images and the input image. To evaluate the impact of feature encoder architecture, we conducted ablation studies with DINOv2 [70] as an alternative feature extractor. For the DINOv2 implementation, we extracted features from the final intermediate layer, resulting in a feature representation $\phi(\cdot)\in\mathbb{R}^{1536\times 16\times 16}$. These features were then used to construct semantic difference maps following the same methodology applied to VGG features.
Qualitative and quantitative comparisons between static camera trajectory (SV3D), adaptive trajectory with DINOv2, and adaptive trajectory with VGG16 are presented in Figure 12 and Table 2, respectively. Each method was evaluated using both NeUS [2] and IM [5] reconstruction pipelines.
Figure 12: Qualitative visual comparisons between 3D reconstruction with static camera trajectory (SV3D), adaptive camera trajectory with DINOv2 (DINO) and adaptive camera trajectory with VGG16 (Ours) on the GSO dataset. Please zoom in for a closer inspection. Meshes are reconstructed with both NeuS and LRM; the latter is noted with the subscript. It can be seen that adaptive camera trajectory provides additional geometric awareness in comparison to static camera trajectory, regardless.
[…]
SV3D(u) [1] 4.93 2.79 18.4
[…]
$\text{Ours}_{\text{DINO}}$ 4.59 3.16 16.9
$\text{Ours}_{\text{DINO}_{\text{IM}}}$ 4.55 3.40 16.4
Ours 4.42 3.52 16.6
$\text{Ours}_{\text{IM}}$ 4.51 3.42 16.2
Table 2: Quantitative results of single-view 3D reconstruction on the full GSO dataset with static camera trajectory and adaptive trajectories with DINOv2 [70] and VGG16 [4] as feature extractors, respectively. Numbers in bold are ranked first.
The results demonstrate that both feature extractors contribute to developing more effective camera paths, yielding improvements in both qualitative and quantitative metrics compared to static trajectories. Notably, VGG16 outperforms DINOv2 in this specific application. We hypothesize that this performance difference stems from VGG16’s superior capability in processing low-level geometric features present in our grayscale slice images, whereas DINOv2 was primarily trained on full-color RGB images and may be less attuned to the specific characteristics of our data.
Appendix F Orbital Camera Trajectories
As described in the main paper, we implement a closed-loop orbital camera trajectory to optimize viewpoint sampling. The camera orbit is structured into four quarters, with camera elevation changing at a constant rate within each quarter. Our path planning algorithm determines the increment steps for the first two segments, while the final two segments mirror these changes with negated values to complete the loop.
This design approach is motivated by two key considerations:
  • Search Space Reduction: Camera pose planning exists in a continuous space with virtually infinite candidate positions. By discretizing elevation changes, we significantly reduce the search space, enabling efficient execution of combinatorial optimization to determine an effective path.
  • Constraint Balance: The closed-loop structure enforces zero total variation in elevation, creating a balance between exploring diverse elevation angles while remaining within the operational capabilities of the video diffusion model.
Our empirical observations indicate that unconstrained greedy algorithms—those without total variation constraints in elevation and closed-loop confinement—frequently produce steeply descending trajectories that result in distorted video generation outputs.
While we acknowledge the current limitations of our approach in exploring asymmetric objects, our method demonstrates significant improvements over static camera trajectories. We believe this work provides valuable insights for future research in reconstruction-centric camera path planning.
Appendix G Complementary Role in Direct 3D supervision
Recent methods employing direct 3D supervision [9, 6] have demonstrated superior performance in 3D generation and reconstruction, achieving high-quality results with faster inference than multiview-based approaches. To investigate this further, we conduct an ablation study on Trellis [9] comparing single image input against multiview inputs sampled from orbital camera trajectories. For multiview evaluation, we sample 6 views on static camera trajectories and 6 views on adaptive camera trajectories, forming three comparison groups: Trellis, Multi-Trellis Static, and Multi-Trellis Adaptive.
The qualitative results in Figure 13 reveal that multiview input Trellis does not consistently outperform its single-view counterpart. This degradation can be attributed to blur and distortions introduced by hallucinated multiview images, which compromise geometric detail quality as evidenced in side-by-side comparisons. However, when input views exhibit significant self-occlusion—as observed in rows 1, 2, and 4 of the right column in Figure 13—additional multiviews enhance geometric awareness in occluded regions, supporting more regularized and occlusion-aware reconstruction. Notably, adaptive multiviews consistently outperform static multiviews in these scenarios due to their superior occlusion-revelation capabilities.
While these advances in direct 3D supervision might suggest that multiview reconstruction is becoming obsolete, we contend that both paradigms offer complementary strengths. Although multiview inputs do not universally improve results, particularly when hallucinated views suffer from distortion or inconsistency, our findings indicate that multiview generation and direct 3D reconstruction can synergistically complement each other when views are strategically selected. We hope this exploration of adaptive trajectory planning will inspire further research into harnessing the combined potential of these two paradigms.
Figure 13: 3D reconstruction results obtained using Trellis [9] with single image input and multiview image inputs sourced from static and adaptive camera trajectories. Single-view inputs consistently produce cleaner results with richer geometric details, as multiview inputs introduce additional blur from hallucinated views, as universally observed across all samples. However, multiview inputs demonstrate advantages for handling self-occlusion scenarios. Notable examples include R1 (Right column, row 1), R2, and R4, where the dog bowl is reconstructed with a solid base, the distorted sheep model is regularized, and the dual-head feature of the dragon is properly emphasized. These cases highlight how strategic multiview selection can enhance reconstruction quality in geometrically challenging scenarios despite the general trade-off in detail fidelity.